how many mL of 0.15 M NaOH are needed to titrate to the end point of 15.5 mL of 0.250 M H2SO4?
1
Expert's answer
2020-11-17T10:35:03-0500

H2SO4 + 2NaOH = Na2SO4 +2H2O

CaVa / CbVb = na/nb

0.250×15.5 / (0.15 × Vb) = 1/2

Vb = 2×0.250×15.5 / 0.15 = 51.67mL